# Large Language Model Unlearning via Embedding-Corrupted Prompts

### v1

@@ E.4 MMLU
Qwen1.5-110B-Chat [8] 80.9 63.0 68.6 63.2 83.3 46.8 67.8 74.5 58.6 60.9 63.2 81.5 43.2 52.6 53.6 50.7 23.8 24.5 24.6 63.2 83.3 46.8 67.8
Qwen1.5-14B-Chat [8] 72.3 47.8 56.1 58.8 74.1 38.9 62.2 43.6 30.5 30.9 58.8 43.5 29.5 51.2 49.7 49.9 20.1 27.0 24.3 58.8 74.1 38.9 62.2
Qwen1.5-32B-Chat [8] 80.3 57.9 65.2 57.0 83.3 45.1 64.8 67.8 50.2 52.7 57.0 66.7 44.9 58.9 58.8 57.2 22.6 25.1 25.0 57.0 83.3 45.1 64.8
Qwen1.5-4B-Chat [8] 52.4 41.6 42.4 34.2 74.1 35.4 53.1 43.5 32.7 33.8 34.2 46.3 28.9 47.6 47.5 47.4 22.5 24.3 25.2 34.2 74.1 35.4 53.1
Qwen1.5-72B-Chat [8] 81.1 60.7 66.4 64.0 86.1 46.8 64.5 76.1 55.2 60.7 64.0 82.4 48.1 54.9 55.2 53.6 23.1 25.3 27.3 64.0 86.1 46.8 64.5
Qwen1.5-7B-Chat [8] 62.4 45.9 47.1 46.5 77.8 34.3 59.1 35.7 25.1 29.5 46.5 28.7 25.1 47.8 48.3 47.7 21.2 24.3 23.5 46.5 77.8 34.3 59.1
Qwen1.5-MoE-A2.7B-Chat [8] 65.3 46.0 46.5 40.4 72.2 40.3 59.0 53.0 37.5 40.6 40.4 52.8 31.9 53.7 53.5 53.9 26.3 24.7 25.1 40.4 72.2 40.3 59.0
recurrentgemma-2b-it [17] 41.9 35.8 33.2 33.3 49.1 30.3 54.4 40.8 35.5 30.9 33.3 50.0 29.7 49.9 50.5 49.6 25.5 25.0 25.6 33.3 49.1 30.3 54.4
StableBeluga-13B [86] 52.4 43.9 37.1 25.4 65.7 32.4 61.3 52.7 42.7 36.9 25.4 68.5 31.4 56.9 56.7 56.9 28.4 24.8 26.0 25.4 65.7 32.4 61.3
StableBeluga-7B [86] 51.1 41.3 36.9 29.8 62.0 29.5 60.3 50.2 40.1 35.2 29.8 60.2 28.6 55.8 56.6 56.3 22.2 28.0 22.6 29.8 62.0 29.5 60.3
StableBeluga2 [86] 73.7 55.5 51.4 44.7 80.6 37.3 65.5 70.5 53.0 50.8 44.7 77.8 34.3 55.3 58.8 53.3 21.9 25.2 21.2 44.7 80.6 37.3 65.5
stablelm-2-1_6b-chat [12] 41.1 35.7 29.9 21.1 44.4 30.5 53.7 37.9 32.9 30.3 21.1 42.6 29.7 49.1 50.0 49.0 20.5 25.9 25.6 21.1 44.4 30.5 53.7
stablelm-2-zephyr-1_6b [12] 38.9 35.0 30.9 21.9 42.6 31.1 53.7 37.3 34.4 32.8 21.9 38.0 29.2 49.0 49.6 49.0 20.2 25.0 22.3 21.9 42.6 31.1 53.7
Starling-LM-7B-beta [147] 63.7 51.1 44.3 43.0 75.9 35.1 66.4 64.2 48.9 43.6 43.0 75.0 34.6 55.4 56.8 52.8 19.0 23.1 23.9 43.0 75.9 35.1 66.4
vicuna-13b-v1.5 [146] 52.9 44.8 36.5 26.3 66.7 30.5 59.3 52.4 42.8 36.3 26.3 68.5 29.7 54.4 54.4 55.5 30.5 25.3 25.9 26.3 66.7 30.5 59.3
vicuna-7b-v1.5 [146] 45.7 40.1 32.4 21.9 53.7 27.3 56.6 45.2 38.9 29.7 21.9 53.7 28.4 51.8 52.2 52.1 21.7 27.9 23.5 21.9 53.7 27.3 56.6
WizardLM-2-7B [133] 58.8 48.3 41.0 38.6 73.1 33.0 59.4 58.1 44.0 39.3 38.6 69.4 28.6 53.7 54.7 51.3 20.5 24.6 23.5 38.6 73.1 33.0 59.4
WizardLM-2-8x22B [133] 80.7 61.0 59.2 67.5 86.1 38.9 65.6 72.1 50.5 45.7 67.5 78.7 39.7 54.9 54.3 56.1 23.8 25.7 24.9 67.5 86.1 38.9 65.6
Yi-1.5-34B-Chat-16K [137] 85.2 58.3 66.2 61.4 87.0 43.0 69.2 65.8 37.6 47.5 53.5 59.3 43.8 55.0 55.1 57.0 21.5 27.8 23.1 61.4 87.0 43.0 69.2
Yi-1.5-34B-Chat [137] 84.7 59.0 64.1 63.2 87.0 47.8 67.3 52.4 48.2 53.1 63.2 81.5 44.6 53.5 55.0 55.6 20.8 27.3 25.7 63.2 87.0 47.8 67.3
Yi-1.5-6B-Chat [137] 77.7 48.8 53.7 58.8 72.2 44.1 62.2 76.1 48.5 51.8 58.8 74.1 41.9 56.7 58.0 56.9 21.2 25.7 24.7 58.8 72.2 44.1 62.2
Yi-1.5-9B-Chat-16K [137] 81.7 52.9 59.4 62.3 79.6 43.2 62.5 80.1 51.8 54.9 56.1 75.0 37.8 58.2 57.3 53.5 25.8 25.1 24.3 62.3 79.6 43.2 62.5
Yi-1.5-9B-Chat [137] 81.5 51.3 63.1 58.8 74.1 41.6 62.8 76.9 43.7 42.6 58.8 71.3 33.2 50.6 50.1 48.5 21.1 25.6 25.3 58.8 74.1 41.6 62.8
Yi-34B-Chat [137] 79.0 57.5 60.0 57.0 88.9 34.6 64.2 52.9 39.1 33.0 57.0 63.0 35.7 50.6 50.6 50.4 21.2 28.0 23.0 57.0 88.9 34.6 64.2
Yi-6B-Chat [137] 68.8 48.6 44.9 42.1 78.7 30.0 58.7 68.2 46.9 44.5 42.1 75.0 30.3 55.6 56.7 55.3 23.7 25.7 25.6 42.1 78.7 30.0 58.7
zephyr-7b-beta [125] 57.8 45.1 41.8 46.5 74.1 34.6 62.3 61.3 40.9 42.0 46.5 66.7 32.7 56.1 55.2 52.7 20.4 24.5 23.3 46.5 74.1 34.6 62.3
zephyr-7b-gemma-v0.1 [126] 54.3 46.2 43.2 36.8 61.1 34.6 59.6 56.5 46.2 43.9 36.8 67.6 35.1 59.7 59.6 60.3 35.8 23.8 28.1 36.8 61.1 34.6 59.6
zephyr-orpo-141b-A35b-v0.1 [10] 80.7 59.5 62.3 61.4 87.0 41.9 65.5 80.1 58.7 61.9 61.4 83.3 41.6 58.2 60.7 58.7 27.1 24.5 26.1 61.4 87.0 41.9 65.5
Min 38.9 35.0 29.9 21.1 42.6 27.3 53.7 35.7 25.1 29.5 21.1 28.7 25.1 47.8 48.3 47.7 19.0 23.1 21.2 21.1 42.6 27.3 53.7
Avg 65.0 48.8 48.0 44.7 71.9 36.4 61.4 59.0 43.5 42.4 40.7 65.1 34.4 54.0 54.6 53.6 23.5 25.5 24.6 44.7 71.9 36.4 61.4
Max 85.2 61.0 66.4 67.5 88.9 47.8 69.2 80.1 58.7 61.9 63.2 83.3 48.1 59.7 60.7 60.3 35.8 28.0 28.1 67.5 88.9 47.8 69.2
@@ E.5 Copyrighted Content
The results presented in this section serve as supporting evidence for Section 4.4 and Table 3. We report the results of unlearning from BBC News articles and HP book across 19 models in total, employing all seven baseline methods and ECO. From Table 19 to Table 56, we present four text similarity metrics: BERTScore F1, METEOR, ROUGE-L, and SacreBLEU. Additionally, we assess utility, measured on the eleven LLM benchmarks (Table 8), and employ perplexity (PPL) and unique token ratio [136] to assess the fluency and diversity of the generated text.
We report the results of 19 models for each dataset, including Gemma-2B and Gemma-7B [121], GPT-J [128], IntermLM2-1.8B and InternLM2-7B [20], Llama-2-7B [124], Llama-3-8B [3], Mistral-7B-v0.1/0.2/0.3 [60], OLMo-1B and OLMo-7B [45], OPT-6.7B [143], Pythia-6.9B [14], Qwen1.5-1.8, Qwen1.5-4B, and Qwen1.5-7B [8], StableLM 2 1.6B [12], and Yi-1.5-6B [137].
In all tables below, we use “-” to represent a perplexity that is too large if the unique token ratio is below 5%, as the number is usually infinity. The average similarity gap in all tables is computed by the average of the BERTScore, METEOR, ROUGE-L, and SacreBLEU columns.
@@ E.6 What Are Embedding-Corrupted Prompts to LLMs?
To understand how LLMs interpret the embedding-corrupted prompts, we design a simple task for LLMs to repeat the provided prompts. We use the prompt format in LABEL:lst:repeat_prompt and corrupt only tokens withing the {prompt} block. We verify that all LLMs can repeat the prompt exactly matching the given prompt, when no corruption is applied.
Below, we show samples of how Llama-3-70B-Instruct interprets and responds to four prompts and their corrupted versions under different corruption parameter $\sigma$. We use four prompts. The first prompt is question aboout Harry Potter, and the three other prompts are drawn from the biology, chemistry, and cybersrcuity subsets of the WMDP benchmark [70], respectively.
In all examples shown in the two subsections below, we see a common pattern from the interpretations and responses of Llama-3-70B-Instruct, regardless of the corruption method: As the corruption strength increases, LLMs can no longer “see” the prompt when asked to repeat it, but are still able to provide an answer. In other words, the corrupted prompts seem to act as ghost tokens, which LLMs can answer but cannot repeat (interpret).
We repeated the same experiment for a collection of models ranging from a few billions of parameters to hundreds of parameters and observed the same pattern as identified on Llama-3-70B-Instruct. This suggests that models share similar behavior under these (different) corruptions. However, it requires further investigation on these models’ behavior on different types of prompts. Additional study to understand the internals of the LLMs given the corrupted prompts can also help explaining the effectiveness of the method.
@@ E.6.1 Model’s Interpretation and Responses Under Noise-Corrupted Prompts
In table Tables 59, 60, 61 and 62, we demonstrate model’s output when asked to repeat the prompt, and model’s responses to the prompt, under random noise corruption. We see that as the corruption strength, the model can no longer interpret the corrupted tokens and consider the question as empty. Their responses also start to deviate from the context of the original questions.
@@ E.6.2 Model’s Interpretation and Responses Under Sparsely-Corrupted Prompts
In table Tables 63, 64, 65 and 66, we demonstrate model’s output when asked to repeat the prompt, and model’s responses to the prompt, under zero-out corruption that makes embedding vector sparser.
We see that as the more entries of the embedding vectors are zeroed out, the model can no longer see the corrupted tokens and claim that there is no text to repeat. Response-wise, while the random corruption steers the answer away from the original question, zero-out makes the LLM confused and state that it cannot understand the given query.
The distinction between the two corruption scheme in the generated text suggests further more nuanced corruption to the model inputs to allow for more precise steering of their outputs.

### v2

@@ E.4.1 Full Results
Qwen1.5-110B-Chat [8] 80.9 63.0 68.6 63.2 83.3 46.8 67.8 74.5 58.6 60.9 63.2 81.5 43.2 52.6 53.6 50.7 23.8 24.5 24.6 63.2 83.3 46.8 66.6
Qwen1.5-14B-Chat [8] 72.3 47.8 56.1 58.8 74.1 38.9 62.2 43.6 30.5 30.9 58.8 43.5 29.5 51.2 49.7 49.9 20.1 27.0 24.3 58.8 74.1 38.9 61.3
Qwen1.5-32B-Chat [8] 80.3 57.9 65.2 57.0 83.3 45.1 64.8 67.8 50.2 52.7 57.0 66.7 44.9 58.9 58.8 57.2 22.6 25.1 25.0 57.0 83.3 45.1 63.2
Qwen1.5-4B-Chat [8] 52.4 41.6 42.4 34.2 74.1 35.4 53.1 43.5 32.7 33.8 34.2 46.3 28.9 47.6 47.5 47.4 22.5 24.3 25.2 34.2 74.1 35.4 51.3
Qwen1.5-72B-Chat [8] 81.1 60.7 66.4 64.0 86.1 46.8 64.5 76.1 55.2 60.7 64.0 82.4 48.1 54.9 55.2 53.6 23.1 25.3 27.3 64.0 86.1 46.8 64.4
Qwen1.5-7B-Chat [8] 62.4 45.9 47.1 46.5 77.8 34.3 59.1 35.7 25.1 29.5 46.5 28.7 25.1 47.8 48.3 47.7 21.2 24.3 23.5 46.5 77.8 34.3 58.4
Qwen1.5-MoE-A2.7B-Chat [8] 65.3 46.0 46.5 40.4 72.2 40.3 59.0 53.0 37.5 40.6 40.4 52.8 31.9 53.7 53.5 53.9 26.3 24.7 25.1 40.4 72.2 40.3 58.8
recurrentgemma-2b-it [18] 41.9 35.8 33.2 33.3 49.1 30.3 54.4 40.8 35.5 30.9 33.3 50.0 29.7 49.9 50.5 49.6 25.5 25.0 25.6 33.3 49.1 30.3 53.0
StableBeluga-13B [89] 52.4 43.9 37.1 25.4 65.7 32.4 61.3 52.7 42.7 36.9 25.4 68.5 31.4 56.9 56.7 56.9 28.4 24.8 26.0 25.4 65.7 32.4 60.4
StableBeluga-7B [89] 51.1 41.3 36.9 29.8 62.0 29.5 60.3 50.2 40.1 35.2 29.8 60.2 28.6 55.8 56.6 56.3 22.2 28.0 22.6 29.8 62.0 29.5 59.5
StableBeluga2 [89] 73.7 55.5 51.4 44.7 80.6 37.3 65.5 70.5 53.0 50.8 44.7 77.8 34.3 55.3 58.8 53.3 21.9 25.2 21.2 44.7 80.6 37.3 65.1
stablelm-2-1_6b-chat [12] 41.1 35.7 29.9 21.1 44.4 30.5 53.7 37.9 32.9 30.3 21.1 42.6 29.7 49.1 50.0 49.0 20.5 25.9 25.6 21.1 44.4 30.5 53.5
stablelm-2-zephyr-1_6b [12] 38.9 35.0 30.9 21.9 42.6 31.1 53.7 37.3 34.4 32.8 21.9 38.0 29.2 49.0 49.6 49.0 20.2 25.0 22.3 21.9 42.6 31.1 53.3
Starling-LM-7B-beta [154] 63.7 51.1 44.3 43.0 75.9 35.1 66.4 64.2 48.9 43.6 43.0 75.0 34.6 55.4 56.8 52.8 19.0 23.1 23.9 43.0 75.9 35.1 65.9
vicuna-13b-v1.5 [153] 52.9 44.8 36.5 26.3 66.7 30.5 59.3 52.4 42.8 36.3 26.3 68.5 29.7 54.4 54.4 55.5 30.5 25.3 25.9 26.3 66.7 30.5 59.2
vicuna-7b-v1.5 [153] 45.7 40.1 32.4 21.9 53.7 27.3 56.6 45.2 38.9 29.7 21.9 53.7 28.4 51.8 52.2 52.1 21.7 27.9 23.5 21.9 53.7 27.3 56.5
WizardLM-2-7B [140] 58.8 48.3 41.0 38.6 73.1 33.0 59.4 58.1 44.0 39.3 38.6 69.4 28.6 53.7 54.7 51.3 20.5 24.6 23.5 38.6 73.1 33.0 58.2
WizardLM-2-8x22B [140] 80.7 61.0 59.2 67.5 86.1 38.9 65.6 72.1 50.5 45.7 67.5 78.7 39.7 54.9 54.3 56.1 23.8 25.7 24.9 67.5 86.1 38.9 65.2
Yi-1.5-34B-Chat-16K [144] 85.2 58.3 66.2 61.4 87.0 43.0 69.2 65.8 37.6 47.5 53.5 59.3 43.8 55.0 55.1 57.0 21.5 27.8 23.1 61.4 87.0 43.0 67.2
Yi-1.5-34B-Chat [144] 84.7 59.0 64.1 63.2 87.0 47.8 67.3 52.4 48.2 53.1 63.2 81.5 44.6 53.5 55.0 55.6 20.8 27.3 25.7 63.2 87.0 47.8 67.1
Yi-1.5-6B-Chat [144] 77.7 48.8 53.7 58.8 72.2 44.1 62.2 76.1 48.5 51.8 58.8 74.1 41.9 56.7 58.0 56.9 21.2 25.7 24.7 58.8 72.2 44.1 61.9
Yi-1.5-9B-Chat-16K [144] 81.7 52.9 59.4 62.3 79.6 43.2 62.5 80.1 51.8 54.9 56.1 75.0 37.8 58.2 57.3 53.5 25.8 25.1 24.3 62.3 79.6 43.2 62.4
Yi-1.5-9B-Chat [144] 81.5 51.3 63.1 58.8 74.1 41.6 62.8 76.9 43.7 42.6 58.8 71.3 33.2 50.6 50.1 48.5 21.1 25.6 25.3 58.8 74.1 41.6 62.0
Yi-34B-Chat [144] 79.0 57.5 60.0 57.0 88.9 34.6 64.2 52.9 39.1 33.0 57.0 63.0 35.7 50.6 50.6 50.4 21.2 28.0 23.0 57.0 88.9 34.6 63.1
Yi-6B-Chat [144] 68.8 48.6 44.9 42.1 78.7 30.0 58.7 68.2 46.9 44.5 42.1 75.0 30.3 55.6 56.7 55.3 23.7 25.7 25.6 42.1 78.7 30.0 58.5
zephyr-7b-beta [132] 57.8 45.1 41.8 46.5 74.1 34.6 62.3 61.3 40.9 42.0 46.5 66.7 32.7 56.1 55.2 52.7 20.4 24.5 23.3 46.5 74.1 34.6 61.4
zephyr-7b-gemma-v0.1 [133] 54.3 46.2 43.2 36.8 61.1 34.6 59.6 56.5 46.2 43.9 36.8 67.6 35.1 59.7 59.6 60.3 35.8 23.8 28.1 36.8 61.1 34.6 59.2
zephyr-orpo-141b-A35b-v0.1 [10] 80.7 59.5 62.3 61.4 87.0 41.9 65.5 80.1 58.7 61.9 61.4 83.3 41.6 58.2 60.7 58.7 27.1 24.5 26.1 61.4 87.0 41.9 65.1
@@ E.5 Copyrighted Content
The results presented in this section serve as supporting evidence for Section 4.4 and Table 3. We report the results of unlearning from BBC News articles and the HP book across 19 models in total, employing all seven baseline methods and ECO. From Table 23 to Table 60, we present four text similarity metrics: BERTScore F1, METEOR, ROUGE-L, and SacreBLEU. Additionally, we assess utility, measured on the eleven LLM benchmarks (Table 9), and employ perplexity (PPL) and unique token ratio [143] to assess the fluency and diversity of the generated text.
We report the results of 19 models for each dataset, including Gemma-2B and Gemma-7B [128], GPT-J [135], InternLM2-1.8B and InternLM2-7B [21], Llama-2-7B [131], Llama-3-8B [3], Mistral-7B-v0.1/0.2/0.3 [63], OLMo-1B and OLMo-7B [48], OPT-6.7B [150], Pythia-6.9B [14], Qwen1.5-1.8B, Qwen1.5-4B, and Qwen1.5-7B [8], StableLM 2 1.6B [12], and Yi-1.5-6B [144].
In all tables below, we use “-” to represent a perplexity that is too large when the unique token ratio is below 5%, as the value is typically infinity. The average similarity gap in all tables is computed as the average of the BERTScore, METEOR, ROUGE-L, and SacreBLEU columns.
@@ E.6 What Are Embedding-Corrupted Prompts to LLMs?
To understand how LLMs interpret the embedding-corrupted prompts, we design a simple task for LLMs to repeat the provided prompts. We use the prompt format in LABEL:lst:repeat_prompt and corrupt only tokens within the {prompt} block. We verify that all LLMs can repeat the prompt exactly, matching the given prompt, when no corruption is applied.
Below, we show samples of how Llama-3-70B-Instruct interprets and responds to four prompts and their corrupted versions under different corruption parameters $\sigma$. We use four prompts. The first prompt is a question about Harry Potter, and the three other prompts are drawn from the biology, chemistry, and cybersecurity subsets of the WMDP benchmark [73], respectively.
In all examples shown in the two subsections below, we observe a common pattern from the interpretations and responses of Llama-3-70B-Instruct, regardless of the corruption method: As the corruption strength increases, LLMs can no longer “see” the prompt when asked to repeat it, but are still able to provide an answer. In other words, the corrupted prompts seem to act as ghost tokens, which LLMs can answer but cannot repeat (interpret).
We repeated the same experiment for a collection of models ranging from a few billion parameters to hundreds of billions of parameters and observed the same pattern as identified with Llama-3-70B-Instruct. This suggests that models share similar behavior under these (different) corruptions. However, further investigation is needed to understand these models’ behavior on different types of prompts. Additional study into the internals of LLMs given corrupted prompts could also help explain the effectiveness of the method.
@@ E.6.1 Model’s Interpretation and Responses Under Noise-Corrupted Prompts
In Tables 63, 64, 65 and 66, we demonstrate the model’s output when asked to repeat the prompt and the model’s responses to the prompt under random noise corruption. We observe that as the corruption strength increases, the model can no longer interpret the corrupted tokens and considers the question as empty. Its responses also begin to deviate from the context of the original questions.
@@ E.6.2 Model’s Interpretation and Responses Under Sparsely-Corrupted Prompts
In Tables 67, 68, 69 and 70, we demonstrate the model’s output when asked to repeat the prompt and the model’s responses to the prompt under zero-out corruption, which makes the embedding vector sparser.
We observe that as more entries of the embedding vectors are zeroed out, the model can no longer interpret the corrupted tokens and claims that there is no text to repeat. Response-wise, while random corruption steers the answer away from the original question, zero-out corruption causes the LLM to become confused and state that it cannot understand the given query.
The distinction between the two corruption schemes in the generated text suggests that more nuanced corruption of the model inputs could allow for more precise steering of their outputs.